OpenStudy (anonymous):

A can holds 3 tennis balls. A box holds 4 cans. A case holds 6 boxes. How many tennis balls are in 10 cases?

OpenStudy (anonymous):

Answer choices please?

OpenStudy (anonymous):

There are no answer choices :(

OpenStudy (anonymous):

Ohh true, okay give me a sec xD I dont wanna mess up.

OpenStudy (anonymous):

Do you need me to show work?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I believe it is 792. What I did was counted out the amount of tennis balls in a box and case. The amount of tennis balls in a case is 72, so I added 72 10 times to get 792. If you want to double check my addition for 72 + 72 ten times, go ahead.

OpenStudy (anonymous):

i think it's 720 im pretty sure

OpenStudy (anonymous):

Ah okay, then it's 720. I must of added more than I needed, because I did actually get 720, but I added more.

OpenStudy (anonymous):

3*4=12 12*6=72 72*10=720

OpenStudy (anonymous):

Thanks a lot though

OpenStudy (anonymous):

Yeah, I added one extra 72 on accident.

OpenStudy (anonymous):

Best of luck to you :)

OpenStudy (anonymous):

Thanks XD